# TD-Orch: Scalable Load-Balancing for Distributed Systems with Applications to Graph Processing

### v1

@@ 4.2. End-to-End Results
(c) PR on Twitter.
Figure 7. Strong scalability on Twitter with 1–16 machines.
4.3. Scalability
Strong Scalability. We run the SSSP, BC, and PR algorithms on the Twitter dataset using a different setup of 1, 2, 4, 8, and 16 machines. The results in Figure 7 demonstrate that TDO-GP exhibits near linear scalability across all machine configurations, while other methods either show poor scalability when increasing the number of machines, or exhibit poor performance. Similar to Table 2, the only case we are not winning is the PR against Graphite, where further details on this behavior will be discussed in Section 4.5.
Weak Scalability. We also evaluate weak scalability on graph systems by varying the number of machines while keeping the graph size per machine fixed. To construct graphs in such scale, we employ two graph generation methods: Erdős-Rényi (ER) (erd6s1960evolution), which produces unskewed graphs, and Barabási–Albert (BA) (barabasi1999emergence), which produces skewed graphs. The average number of edges per machine is fixed at 40M for a fair comparison, where in the 16-machine case the graph is in similar size as uk-2005. For the BA model, we set the power-law exponent to $\gamma=2.2$, consistent with the measured skew observed in natural graphs reported by PowerGraph (gonzalez2012powergraph). We run both PR and BC on $P=1,2,4,8,16$ machines. In Figure 9, the x-axis represents the number of machines and the y-axis represents the execution time. The results show that TDO-GP maintains nearly constant runtime across different machine counts, demonstrating excellent weak scalability, whereas prior works exhibit significant scalability degradation.
(a) ER-PR
(b) ER-BC
(c) BA-PR
(d) BA-BC
Figure 8. Weak scalability results. $y$-axis is runtime (in seconds) and $x$-axis is the number of machines used.
4.4. Ablation Study
Execution Time Breakdown. As shown in Figure 9, on 16 machines the time spent on pure local computation is effectively reduced across all algorithms. At the same time, the communication time (mainly for propagating vertex values and the cost of synchronization rounds) and the overhead time (including data preparation and data receiving for multi-machine transmission) are also non-negligible, and together contribute significantly to the overall execution time.
Figure 9. Runtime on Twitter with 16 machines. Breakdown into Communication, Computation, and Overhead.
Sensitivity to Each Optimization. We first demonstrate the effectiveness of TD-Orch within TDO-GP by comparing its final design against a baseline of Ligra with direct-pull, which performs no load balancing across machines. As shown in Table 3, our BC results on the Twitter graph achieve up to $220\times$ speedup.
Additionally, we evaluate the sensitivity of TDO-GP to the implementation techniques introduced in Section 3.2. Using a controlled-variable approach, we measure the slowdown incurred when removing each optimization family from the fully optimized TDO-GP. Table 4 reports the results on the Twitter dataset using 4, 8, and 16 machines. The results show that disabling any single optimization family causes substantial performance degradation, with slowdowns of up to $5.72\times$, underscoring the critical importance of each technique in achieving high practical efficiency.
Table 3. Runtime (in seconds) of BC on Twitter.
Machine Num. 1 4 8 16
Ligra-Dist 5.36 563.11 670.71 321.60
@@ 4.4. Ablation Study
Table 4. Slowdown when removing individual techniques (see Section 3.2) from the fully optimized TDO-GP.
@@
Removing 3.2.1: Optimized Global Communication.
@@
Removing 3.2.2: Work-Efficient Local Computation.
@@
Removing 3.2.3: Aligned Coordination.
@@
4.5. Ablation Study on NUMA Effects
As discussed in Sections 4.2 and 4.3, the only case where TDO-GP does not outperform prior systems is PageRank (PR), and only against Graphite (mofrad2020graphite). TDO-GP is built upon ParlayLib (blelloch2020parlaylib), a fast parallel library that achieves strong scalability with respect to the total number of cores, but is NUMA-oblivious. While this design works well on machines with high-speed NUMA interconnects, our 16-machine cluster has budget constraints: each machine contains four NUMA nodes connected in a square topology, where certain NUMA-to-NUMA accesses are significantly slower. This architectural limitation hinders the performance of TDO-GP on PR.
We conduct two additional ablation studies to evaluate the effectiveness of TDO-GP when NUMA-related limitations are mitigated. First, we perform a multi-machine experiment on the same cluster used in Section 4.2, but restrict execution to a single NUMA node per machine. The results, shown in Table 5, demonstrate that the performance gap between TDO-GP and baselines is substantially reduced. Moreover, TDO-GP achieves a significantly smaller memory footprint compared to Graphite. The remaining minor performance difference arises because ParlayLib (blelloch2020parlaylib), the underlying parallel library of TDO-GP, is optimized for many-core settings, making the single-NUMA configuration less favorable.
We further conduct a single-machine experiment on a more powerful server with superior NUMA interconnects. This machine is equipped with four Intel(R) Xeon(R) E7-8867 v4 @2.40GHz processors, providing a total of 144 cores and 180MB of L3 cache. Unlike our 16-machine cluster, it features an all-to-all interconnection across its four NUMA nodes, eliminating the bottlenecks of square-topology interconnects. The results on the Twitter dataset are presented in Table 6, where TDO-GP consistently outperforms all other baselines by a significant margin. Notably, TDO-GP even surpasses GBBS (dhulipala18scalable), the state-of-the-art single-machine implementation, owing to our optimized lightweight local EdgeMap.
Given the results of these ablation studies, we are confident that TDO-GP can be efficiently deployed to address all the graph problems in this paper on clusters with fast NUMA interconnections, which are typical in high-performance computing environments and modern datacenters.
Table 5. Runtime (in seconds) of PR on Twitter. Each machine uses only 11 NUMA node. OOM is out-of-memory.
@@ 4.5. Ablation Study on NUMA Effects
Table 6. Twitter runtime (in seconds) on Xeon E7-8867 v4.
@@
5. Conclusion
In this paper, we introduce an abstraction of the task–data orchestration problem, unifying a range of distributed applications. We design TD-Orch to realize this orchestration with scalable load balance and low communication. Building on TD-Orch, we develop TDO-GP, a distributed graph processing system enhanced with implementation techniques for graph workloads. Both TD-Orch and TDO-GP are theoretically grounded and achieve high practical performance.
Acknowledgements.
@@ Appendix A Models for Distributed Computing
In this paper, we model and analyze algorithms using the classic Bulk-Synchronous Parallel (BSP) model (valiant1990bridging; valiant1990general; valiant1988optimally). The BSP consists of $P$ machines, each with a sequential processing element (PE) and a local memory. There is no shared global memory. All machines are connected by a point-to-point communication network. All remote accesses to data on other machines are carried out via messages sent over this network. The system operates in bulk-synchronous rounds (supersteps), where periodic barrier synchronizations separate periods of asynchronous computation and communication.
@@
Cost Metrics. The BSP model defines the notion of an $h$-relation of a superstep, which combines the maximum computation work, maximum communication amount and synchronization time over all machines. Due to using maximum when defining $h$-relations, load balance is critical for achieving good performance. In all our analysis, for clarity, we decouple the computation time and communication time, and do not consider the BSP parameters $g$ and $L$ in (valiant1990bridging; gerbessiotis1994direct). Computation time $t$ and communication time $h$ can be combined into a formal BSP result $\Theta(gh+t)$ as long as the synchronization overhead can be hidden by system communication, i.e., $(gh+t)\in\Omega(L)$.
@@
Other Distributed Models. The Massively Parallel Model of Computation (MPC) (im2023massively; karloff2010model; goodrich2011sorting; beame2013communication; andoni2014parallel) family of models regards synchronization as the main bottleneck in distributed systems and takes communication rounds as the only optimization target. We choose the the BSP model because it differs from MPC in two ways: (i) The the BSP model accounts for the total communication among machines rather than ignoring the amount of communication in each round. (ii) The the BSP model also considers computation costs neglected by the MPC model, which turn out to be critical for real-world system designs on emerging distributed hardware (kang2022pimtreevldb; bindschaedler2020hailstorm) other than MapReduce systems  (dean2008mapreduce).
@@
LOCAL/CONGEST models (linial1987distributive; linial1992locality) are a popular model family for distributed systems. Like the models we consider, each machine has only a local view of the computation. These models consider network topology, which we ignore in the BSP model. However, they have similar weaknesses to the MPC model with regards to their metrics: (i) each machine can send/receive up to its maximum allowed communication at each round, whereas we seek to minimize communication, and (ii) computation costs are not considered.
@@
The Processing-in-Memory (PIM) model (kang2021processing) is a recently proposed model for emerging compute-in-memory hardware (mutlu2023primer). The model has a powerful Host CPU and $P$ PIM nodes, each comprised of a PE and a memory. It operates in bulk-synchronous rounds, but unlike the BSP model, all communication must be sent by or to the Host CPU. Furthermore, the model decouples the computation time of the CPU from other processors, resulting in a lack of analysis of load-balance between them.
Appendix B Proof for Theorem 2.1
@@
We will use the balls-into-bins game of Lemma B.1 to prove load balance.
@@
Lemma B.1 (Weighted Balls into Bins  (sanders1996competitive; nievergelt1973binary)).
@@
We set $C=\Theta(1)$ and $F=\Theta(\frac{n}{P\log^{2}{P}})$. (i) Substage 1 takes at most $O(\log_{F}{P})$ rounds and in each round at most $O(n)$ words of messages are sent. Since meta-task sizes are bounded by $O(C\log_{C}{n})\leq{O}(\log{P})$, the weight limit in Lemma B.1 is $O(FC\log_{C}{n})\leq{O}(\frac{n}{P\log{P}})$ and the communication is thus load-balanced. When $n=\Omega(\frac{P\log^{3}{P}}{\log\log{P}})$, $\Theta(\log{F})=\Theta(\log{\frac{n}{P}})$, so the communication time is $O(\frac{n}{P}\log_{\frac{n}{P}}{P})$ in substage 1. Substage 2 takes less communication time than substage 1 because it only requires partial communication over substage 1’s trace, so (i) is proved. (ii) Since meta-tasks have size $O(C\log_{C}{n})$, Lemma B.1 can also be applied. ∎
Appendix C Detailed Orchestration Result
Table 7. Runtime (in seconds) of a weak-scaling microbenchmark of hash table workloads with different numbers of machines under different Zipf parameters $\gamma$.
Workload 2 4 8 16
$\gamma=1.5$
Direct Push 0.1093 0.1747 0.4696 0.5869
Direct Pull 0.0933 0.1572 0.2818 0.3542
Sorting 0.2545 0.3044 0.5397 0.5365
TD-Orch 0.0895 0.1082 0.1590 0.1863
$\gamma=2.0$
Direct Push 0.1697 0.2770 0.5072 0.9655
Direct Pull 0.0903 0.1609 0.2536 0.4579
Sorting 0.4499 0.3080 0.3717 0.4349
TD-Orch 0.0847 0.1280 0.1468 0.1392
$\gamma=2.5$
Direct Push 0.1480 0.3024 0.8190 1.5284
Direct Pull 0.0939 0.1197 0.3183 0.5991
Sorting 0.2030 0.2544 0.5029 0.5771
TD-Orch 0.0936 0.1174 0.2417 0.2937
Appendix D Algorithm for Betweenness Centrality
@@
Algorithm 3 describes the betweenness centrality (BC) computation from a single root vertex, which is commonly used in performance tests in prior works (ShunB2013; dhulipala18scalable; zhu2016gemini).
@@ 35    BCValues[$i$] = BCValues[$i$] / NumPaths[$i$] - 1
Appendix E Details for Implementation Techniques
We present here the implementation techniques used in our graph processing system, which efficiently exploit the data layout and execution flow provided by TD-Orch, leading to the strong performance demonstrated in Section 4.
E.1. Optimized Global Communication
As we will show in Section 4.4, the critical bottleneck in a distributed graph system is the inter-machine communication time. Optimizing global communication is therefore essential. First, we apply deduplication during message passing along the meta-task trees, aggressively merging all repeated items. Second, in dense mode, we reduce broadcast traffic by sending the values of each vertex only to the machines that store its corresponding edges. The mapping of vertices to machines is established during the preprocessing phase.
E.2. Work-Efficient Local Computation
For all local computations in our implementation, we employ work-efficient designs that are both theoretically optimal and practically fast. Our design leverages ParlayLib (blelloch2020parlaylib), a state-of-the-art C++ parallel library providing a wide range of work-efficient parallel algorithms as primitives.
The DistVertexSubset structure consists of a collection of VertexSubset instances, one per machine, where each VertexSubset can independently switch between two representation formats introduced in Ligra (ShunB2013). We introduce two implementation improvements. (i) In the sparse representation, we replace their array of vertices with a phase-concurrent hash table with adaptive sizing, enabling more efficient lookups within VertexSubset. (ii) In the dense representation, we replace the original parallel C++ Boolean-map with a concurrent bitmap, improving cache efficiency.
For local edges whose source and destination vertices reside on the same machine, we bypass the global orchestration framework and instead execute a local EdgeMap. We implement a lightweight EdgeMap entirely using ParlayLib primitives, which, as we show in Section 4.5, is more efficient than the ones in Ligra/GBBS for EdgeMap workloads.
E.3. Aligned Coordination between Components
For better alignment with the parallel thread scheduler in ParlayLib, we avoid using nested ParallelFor, which is commonly seen in graph algorithms but inefficient under such schedulers. Instead, we adopt Flattern, assisted by delayed structures (blelloch2020parlaylib), to transform two-dimensional nested parallelism into one-dimensional parallelism.

### v2

@@ 6.3 Scalability
Figure 7: Strong scaling results on Twitter dataset with 1–16 machines. TDO-GP is significantly faster and more scalable.
(a) PR on ER Graph
(b) PR on BA Graph
(c) BC on ER Graph
(d) BC on BA Graph
Figure 8: Weak scaling results. $y$-axis is runtime (in seconds) and $x$-axis is the number of machines used. TDO-GP exhibits excellent weak-scaling performance and is significantly faster.
6.4 Ablation Study
Execution Time Breakdown. As shown in Figure˜9, on 16 machines, the time spent on pure local computation is effectively reduced across all algorithms. At the same time, the communication time (mainly for propagating vertex values and the cost of synchronization rounds) and the overhead time (including data preparation and data receiving for multi-machine transmission) are also non-negligible, and together contribute significantly to the overall execution time.
Sensitivity to Each Optimization. We first demonstrate the effectiveness of TD-Orch within TDO-GP by comparing its final design against its prototype of Ligra plus direct-pull, which does not use TD-Orch. As shown in Table˜3, our BC results on the Twitter graph achieve up to $220\times$ speedup.
Additionally, we evaluate the sensitivity of TDO-GP to the implementation techniques introduced in §5.2. Using a controlled-variable approach, we measure the slowdown incurred when removing each optimization family from the fully optimized TDO-GP. Table˜4 reports the results on Twitter with $P=4,8,16$, showing that disabling any single optimization family causes substantial performance degradation, with slowdowns of up to $5.72\times$, illustrating the importance of each technique in achieving high practical efficiency.
Figure 9: Runtime on Twitter with 16 machines. Breakdown into Communication, Computation, and Overhead.
Table 3: Runtime (in seconds) of BC on Twitter shows that TD-Orch is a critical building block in TDO-GP.
Number of Machines 1 4 8 16
Ligra-Dist (no TD-Orch) 5.36 563.11 670.71 321.60
@@ 6.4 Ablation Study
Table 4: Slowdown when removing individual techniques (see §5.2) from the fully optimized TDO-GP.
@@
Removing (T1): Optimized Global Communication.
@@
Removing (T2): Work-Efficient Local Computation.
@@
Removing (T3): Aligned Coordination.
@@
6.5 Ablation on PageRank and NUMA Effects
As shown in Table˜2, the only 2 out of the 30 cases where TDO-GP underperforms prior systems are both PageRank (PR), and only against one prior baseline. TDO-GP is built upon ParlayLib [blelloch2020parlaylib], a fast parallel library that achieves good scalability with respect to the total number of cores, but is NUMA-oblivious. While this design works well on machines with high-speed NUMA interconnects, our 16-machine cluster has budget constraints: each machine contains four NUMA nodes connected in a square topology, where certain NUMA-to-NUMA accesses are significantly slower. This architectural limitation hinders the performance of TDO-GP on PR.
We conduct two additional ablation studies to evaluate the effectiveness of TDO-GP when NUMA-related limitations are mitigated. First, we perform a multi-machine experiment on the same cluster used in §6.2, but restrict execution to a single NUMA node per machine. The results, shown in Table˜5, demonstrate that the performance gap between TDO-GP and baselines is substantially reduced. Moreover, TDO-GP has a significantly smaller memory footprint than Graphite. The remaining minor performance difference arises because ParlayLib [blelloch2020parlaylib], the underlying parallel library of TDO-GP, is optimized for many-core settings, making the single-NUMA configuration less favorable.
We further conduct a single-machine experiment on a more powerful server with superior NUMA interconnects. This machine is equipped with four Intel(R) Xeon(R) E7-8867 v4 @2.40GHz processors, providing a total of 144 cores and 180MB of L3 cache. Unlike our 16-machine cluster, it features an all-to-all interconnection across its four NUMA nodes, eliminating the bottlenecks of square-topology interconnects. The results on the Twitter dataset are presented in Table˜6, where TDO-GP consistently outperforms all other baselines by a significant margin. Notably, TDO-GP even surpasses GBBS [dhulipala18scalable], the state-of-the-art single-machine implementation, owing to our optimized lightweight local EdgeMap.
Given the results of these ablation studies, we are confident that TDO-GP can be efficiently deployed to address all the graph problems mentioned in this paper on clusters with fast NUMA interconnections, which are typical in high-performance computing environments and datacenters.
Table 5: Runtime (in seconds) of PR on Twitter. Each machine uses only 11 NUMA node. OOM is out-of-memory.
@@ 6.5 Ablation on PageRank and NUMA Effects
Table 6: Twitter runtime (in seconds) on Xeon E7-8867 v4.
@@
7 Conclusion
In this paper, we introduce an abstraction of the task-data orchestration interface, unifying a wide range of distributed applications. We design TD-Orch to realize this orchestration with scalable load balance and low communication. Building on TD-Orch, we develop TDO-GP, a distributed graph processing system enhanced with implementation techniques for graph workloads. Both TD-Orch and TDO-GP are theoretically grounded and achieve high practical performance.
Acknowledgements
@@ Appendix A Models for Distributed Computing
In this paper, we model and analyze algorithms using the classic Bulk-Synchronous Parallel (BSP) model [valiant1990bridging, valiant1990general, valiant1988optimally]. The BSP consists of $P$ machines, each with a sequential processing element (PE) and a local memory. There is no shared global memory. All machines are connected by a point-to-point communication network. All remote accesses to data on other machines are carried out via messages sent over this network. The system operates in bulk-synchronous rounds (supersteps), where periodic barrier synchronizations separate periods of asynchronous computation and communication.
@@
Cost Metrics. The BSP model defines the notion of an $h$-relation of a superstep, which combines the maximum computation work, maximum communication amount and synchronization time over all machines. Due to using maximum when defining $h$-relations, load balance is critical for achieving good performance. In all our analysis, for clarity, we decouple the computation time and communication time, and do not consider the BSP parameters $g$ and $L$ in [valiant1990bridging, gerbessiotis1994direct]. Computation time $t$ and communication time $h$ can be combined into a formal BSP result $\Theta(gh+t)$ as long as the synchronization overhead can be hidden by system communication, i.e., $(gh+t)\in\Omega(L)$.
@@
Other Distributed Models. The Massively Parallel Model of Computation (MPC) [im2023massively, karloff2010model, goodrich2011sorting, beame2013communication, andoni2014parallel] family of models regards synchronization as the main bottleneck in distributed systems and takes communication rounds as the only optimization target. We choose the the BSP model because it differs from MPC in two ways: (i) The the BSP model accounts for the total communication among machines rather than ignoring the amount of communication in each round. (ii) The the BSP model also considers computation costs neglected by the MPC model, which turn out to be critical for real-world system designs on emerging distributed hardware [kang2022pimtreevldb, bindschaedler2020hailstorm] other than MapReduce systems  [dean2008mapreduce].
@@
LOCAL/CONGEST models [linial1987distributive, linial1992locality] are a popular model family for distributed systems. Like the models we consider, each machine has only a local view of the computation. These models consider network topology, which we ignore in the BSP model. However, they have similar weaknesses to the MPC model with regards to their metrics: (i) each machine can send/receive up to its maximum allowed communication at each round, whereas we seek to minimize communication, and (ii) computation costs are not considered.
@@
The Processing-in-Memory (PIM) model [kang2021processing] is a recently proposed model for emerging compute-in-memory hardware [mutlu2023primer]. The model has a powerful Host CPU and $P$ PIM nodes, each comprised of a PE and a memory. It operates in bulk-synchronous rounds, but unlike the BSP model, all communication must be sent by or to the Host CPU. Furthermore, the model decouples the computation time of the CPU from other processors, resulting in a lack of analysis of load-balance between them.
Appendix B Proof for Theorem˜1
@@
We will use the balls-into-bins game of Lemma˜1 to prove load balance.
@@
Lemma 1 (Weighted Balls into Bins  [sanders1996competitive, nievergelt1973binary]).
@@
When $n=\Omega(P^{\epsilon})$ for some $\epsilon>1$, things become trivial by setting $F=\Theta(n/P)$. Here we only analyze the hard case when $n=O(P\log^{\epsilon}{P})$ for some $\epsilon>1$. We set $C=\Theta(B/\sigma)$ and $F=\Theta(\log{P}/\log{\log{P}})$. (i) Phase 1 takes at most $O(\log_{F}{P})$ rounds and in each round at most $O(n)$ words of messages are sent. Since meta-task sizes are bounded by $O(C\log_{C}{n})\leq{O}\left(\frac{n}{P\log{P}}\right)$, the weight limit in Lemma˜1 is met and the communication is thus load-balanced. When $n=\Omega\left(\frac{PB\log^{3}{P}}{\log\log{P}}\right)$, the communication time is $O(\frac{n}{P}\log_{F}{P})=O(\frac{n}{P}\log_{\frac{n}{P}}{P})$ in phase 1. Phase 2 takes less communication time than phase 1 because it only requires partial communication over phase 1’s trace, so (i) is proved. (ii) Since meta-tasks have size $O(C\log_{C}{n})$, Lemma˜1 can also be applied. ∎
Appendix C Example Algorithms for Graph Problems
Algorithm˜1 presents pseudocode for breadth-first search (BFS), offering a more efficient implementation via DistEdgeMap than directly using the orchestration interface.
@@ Appendix C Example Algorithms for Graph Problems
In addition, Algorithm˜3 is presented here to describe the pseudocodes for the betweenness centrality (BC) computation from a single root vertex, which is commonly used in performance tests in prior works [ShunB2013, dhulipala18scalable, zhu2016gemini].
We also provide the actual C++ implementation of the BC algorithm in Figure˜10 using the DistEdgeMap interface. Despite the complexity of the original algorithm, the code remains highly concise, requiring fewer than 70 lines.
Appendix D Details of Implementation Techniques
We present here the implementation techniques used in our graph processing system, which efficiently exploit the data layout and execution flow provided by TD-Orch, leading to the strong performance demonstrated in §6.
D.1 Optimized Global Communication
As we will show in §6.4, the critical bottleneck in a distributed graph system is the inter-machine communication time. Optimizing global communication is therefore essential. First, we apply deduplication during message passing along the meta-task trees, aggressively merging all repeated items. Second, in dense mode, we reduce broadcast traffic by sending the values of each vertex only to the machines that store its corresponding edges. The mapping of vertices to machines is established during the preprocessing phase.
D.2 Work-Efficient Local Computation
For all local computations in our implementation, we employ work-efficient designs that are both theoretically optimal and practically fast. Our design leverages ParlayLib [blelloch2020parlaylib], a state-of-the-art C++ parallel library providing a wide range of work-efficient parallel algorithms as primitives.
The DistVertexSubset structure consists of a collection of VertexSubset instances, one per machine, where each VertexSubset can independently switch between two representation formats introduced in Ligra [ShunB2013]. We introduce two implementation improvements. (i) In the sparse representation, we replace their array of vertices with a phase-concurrent hash table with adaptive sizing, enabling more efficient lookups within VertexSubset. (ii) In the dense representation, we replace the original parallel C++ Boolean-map with a concurrent bitmap, improving cache efficiency.
For local edges whose source and destination vertices reside on the same machine, we bypass the global orchestration framework and instead execute a local EdgeMap. We implement a lightweight EdgeMap entirely using ParlayLib primitives, which, as we show in §6.5, is more efficient than the ones in Ligra/GBBS for EdgeMap workloads.
D.3 Aligned Inter-Component Coordination
For better alignment with the parallel thread scheduler in ParlayLib, we avoid using nested ParallelFor, which is commonly seen in graph algorithms but inefficient under such schedulers. Instead, we adopt Flattern, assisted by delayed structures [blelloch2020parlaylib], to transform two-dimensional nested parallelism into one-dimensional parallelism.
Input: $G=(V,E)$, where $|V|=n$; StartVertex
Output: Dist[$1:n$]: Distrib. vertex value array
[…]
2 round=1=1
3 Function f ( (u,v): Edge ) :
4  return round IF (Dist[$u$] == round - 1) ELSE -1
5Function write_back ( v: Vertex, aggr_value ) :
6  If Dist[$v$] == -1 and aggr_value != -1 then
7     Dist[$v$] = aggr_value
8     return True
9 return False
10Function BFS ( $G$, StartVertex ) :
11  DistVertexSubset frontier = {StartVertex}
12  Dist[StartVertex] = 0
13  While frontier is not empty do
14     frontier = DistEdgeMap ($G$, frontier, f, write_back, MAX)
15     round += 11
16 return Dist[$1:n$]
Algorithm 2 Breadth-First Search (BFS)
@@ 35    BCValues[$i$] = BCValues[$i$] / NumPaths[$i$] - 1
Figure 10: The actual C++ implementation of betweenness centrality (BC) using our DistEdgeMap interface is highly concise—fewer than 70 lines of code.